OpenStudy (pulsified333):

Assume that the number of defective basketballs produced is related by a linear equation to the total number produced. Suppose that 6 defective balls are produced in a lot of 350, and 15 defective balls are produced in a lot of 475. Find the number of defective balls produced in a lot of 725 balls.

OpenStudy (pulsified333):

@ganeshie8

OpenStudy (pulsified333):

@dan815

OpenStudy (pulsified333):

@jayzdd

OpenStudy (ali2x2):

First find what the linear equation

OpenStudy (ali2x2):

which is

OpenStudy (ali2x2):

y = mx + b m = (525 - 400)/(18 - 9) m = 125/9

OpenStudy (ali2x2):

ok now find b,

OpenStudy (ali2x2):

by plugging in one of the points. Y will be the number of balls produced. The equation can also be set up using y as the number of defective balls. That way your m and b will be different.

OpenStudy (ali2x2):

y = (125/9)x + b 400 = (125/9)(9) + b b = 275 So the equation is y = (125/9)x + 275

OpenStudy (ali2x2):

Now plug in 775 for y

OpenStudy (ali2x2):

775 = (125/9)x + 275 x = 36

OpenStudy (ali2x2):

ok now when 775 balls are produced 36 of em will be defective.

OpenStudy (ali2x2):

Did i help?

OpenStudy (pulsified333):

its for 725 balls not 775

OpenStudy (pulsified333):

@ali2x2

OpenStudy (ali2x2):

?